Substition method problem:
2x-5y=25
X-2y=10

Answer by mananth(16946)

x-2y =10 .............1
y=10-x ---------2
Substitute y in (2)
2x-5y=25
2x-5(10 -x)= 25
2x-50+5x= 25
2x+5x =75
7x=75
/ 7
x= 75/ 7
Plug the value of x in (1)
2x-5y = 25
2*75/ 7-5y=25
150/ 7 + -5y = 25
-5 y = 25/ 7
/ -5
y= - 5/ 7
m.ananth@hotmail.ca